OpenStudy (anonymous):

Which answer shows this equation in standard form? 6 – 2(x – y) = –3x + 5 A. x + 2y = –1 B. x + 2y = 1 C. x – 2y = 1 D. –5x + 2y = –1

OpenStudy (anonymous):

Plz help Medal and Fan to best thing

OpenStudy (anonymous):

@lovielove456767890

OpenStudy (anonymous):

okay I am hear

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

A. because I had the question. 100%right

OpenStudy (tkhunny):

Have you considered PUTTING it in Standard Form? 6 -2(x – y) = –3x + 5 Distributive Property on the left 6 -2x + 2y = –3x + 5 Add 3x 6 + x + 2y = 5 Now what?

OpenStudy (anonymous):

ok can you help with more?

OpenStudy (anonymous):

who the @tkhunny or @lovielove456767890

OpenStudy (anonymous):

@lovielove456767890

OpenStudy (anonymous):

what is the question, and I gave you a medal

OpenStudy (tkhunny):

You have shown neither work nor understanding. I left you one tiny step. Do , please.

OpenStudy (anonymous):

What is the x-intercept of the line with the equation http://static.k12.com/bank_packages/files/media/mathml_821fedbbac96c7e59cd7c82fc9b94616ebef6886_1.gif ( , 0)

OpenStudy (anonymous):

are you from K12

OpenStudy (tkhunny):

Did we solve the last one? Post new questions on new threads, please and NEVER just post a question with no work shown.

OpenStudy (anonymous):

Teen I explained to you how to find the x - intercept on your earlier post, it is the same thing here, just with different numbers, to find it, set y = 0 and solve for x.

OpenStudy (anonymous):

I got confused on this one I wrote down all the thing you told me but when I got to this one I got confused HONEST

OpenStudy (anonymous):

No worries, can you show me where you had troubles, and I can point you towards the right direction.

OpenStudy (anonymous):

Bye, since nobody wants my help anymore

OpenStudy (tkhunny):

So? Show your work and get unconfused. Did you have trouble with the distributive property? Yes or no? Any trouble adding 3x to both sides? Yes or no? Can you subtract constants from both sides? Yes or no? Do you know what "Standard Form" is? Yes or no?

OpenStudy (anonymous):

tkhunny what the heck I am trying to do work NOT argue with people

OpenStudy (anonymous):

yeah not to argue, nobody wants my help anymore. I gave you a medal @TeenWolfGirl because you are nice and my friend. By the way nobody is nice to me in this site